The moment generating fucntion of a random variable X is given by M(t)=(1-t/2)-3.

Find the mean and variance of X


1
Expert's answer
2021-08-09T15:58:15-0400
"mean=\mu=E(X)=M'_t|_{t=0}"

"=(-3(-\dfrac{1}{2})(1-\dfrac{t}{2})^{-4})|_{t=0}=\dfrac{3}{2}"

"E(X^2)=M''_t|_{t=0}"

"=(-4(\dfrac{3}{2})(-\dfrac{1}{2})(1-\dfrac{t}{2})^{-5})|_{t=0}=3"

"Var(X)=E(X^2)-(E(X))^2=3-(\dfrac{3}{2})^2 =\dfrac{3}{4}"